# Spectra of Magnetoroton and Chiral Graviton Modes of Fractional Chern Insulator

### v2

@@ Abstract
Employing the state-of-the-art time-dependent variational principle algorithm, we compute the spectra of charge neutral excitations in the $\nu=1/2$ fractional Chern insulator (FCI) on the Haldane honeycomb lattice model with hard-core bosons. The magnetoroton visualized from the dynamic density structure factor, acquire a minimum gap at finite momentum that can go soft with increasing interaction and give rise to a charge density wave (CDW) at the same wavevector. The sharper and more concentrated spectral weight of the roton mode is observed when approaching the FCI-CDW transition while the single-particle gap remains intact. We also resolve the graviton mode around $\Gamma$ point of the Brillouin zone by measuring the dynamic quadrupolar density correlations. Interestingly, we not only reveal the graviton response is chiral for a certain FCI, but also show the different chiralities of the graviton of opposite-sign Hall conductance for the first time. Our results offer the clear spectral observations of magnetoroton and chiral graviton in a lattice model of FCI and will have relevance towards the on-going experiments of tunable FCIs in quantum moiré materials.
Introduction.— The fractional quantum Hall effect (FQHE) is one of the most exotic macroscopic quantum phenomena caused by the interplay of electron correlation and topology in modern condensed matter physics [1, 2], where 2D electron gas under strong mangetic field form an incompressible quantum fluid with fractional charge and statistics. The lattice analog of the FQHE is called the fractional Chern insulator (FCI) [3, 4, 5, 6], and the zero-magnetic-field FCIs are also known as the fractional quantum anomalous Hall (FQAH) states[4, 7, 8]. The recent breakthroughs in the observations of FQAH states in twisted transition metal dichalocogenide MoTe2 bilayers [9, 10, 11, 12] and in rhombohedral pentalayer graphene/hBN moiré superlattices [13], have attracted broad attentions and the discussions of the properties of these FQAH states and their transitions to charge density wave (CDW) state, metallic state and various anomalous Hall crystal states are actively undertaking [14, 15, 16, 17, 18, 19].
FCIs have been studies in many aspects in connection with FQHE, including the similarity and difference between topological flat bands and Landau levels [20, 21, 22, 23], the correspondence between their wave functions [24] and the ground state entanglement and generalized Pauli principle [6]. As for the collective excitations, the charge-neutral excitations are known to play a dominate role in controlling the properties of the FQHE/FCI systems. For example, the softening of magnetoroton (minimum) is an intrinsic mechanism to trigger phase transitions from FQHE/FCI to CDW [25, 26, 27, 28, 29, 30, 4, 31, 32, 33, 34]. The magnetoroton in the long-wavelength limit is known as the chiral graviton [35], reflecting the oscillation of the geometrical properties, which carries the spin angular momentum $2$ [36, 37, 38, 39, 40, 41]. This geometrical (apart from the topological) description of the FQHE was proposed by Haldane [42] and the experimental observation of chiral gravitons of FQH states was recently reported [43]. Similar as the roton minimum, the graviton mode could also approach the ground-state energy which leads to the FQH nematic state with anisotropic transport [44, 45, 46, 47] and describe quench dynamics [48, 49].
@@
Figure 1: Static density structure factor $S^{A}(\mathbf{k})$ and charge order parameter $\delta^{A}(\mathbf{k})$. Along the parameter path we choose, the system undergoes a sequence of FCI-Solid I-Solid II phase transitions. (a), (b) are the $S^{A}(\mathbf{k})$ of FCI phase ($V_{1}=4,V_{2}=2.5$) and Solid I phase ($V_{1}=4,V_{2}=3.2$). (c), (d) are the real and reciprocal spaces of the honeycomb lattice FCI model in Eq. (1). In (c), $\mathbf{r_{1}}=(\sqrt{3},0),\mathbf{r_{2}}=(-\sqrt{3}/2,3/2)$ are the primitive lattice vectors. The NN, NNN, and NNNN hoppings are depicted with black, blue and green colours, respectively. In (d), the solid hexagon is the first Brillouin zone (BZ) of the model, with high symmetry points labelled. The four paths that are discrete in $\mathbf{b_{2}}$ direction with step $\mathbf{b_{2}}/4$ are where we compute the spectra under the cylinder geometry in DMRG calculation. (e), (f) are the $\delta^{A}(\mathbf{k})$ of Solid I phase ($V_{1}=4,V_{2}=3.2$) and Solid II phase ($V_{1}=4,V_{2}=4.6$). The roton mode locates at $\mathbf{b_{1}}/2$ in (a) condensates and leads to charge order in (e), while similar $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ roton in (b) condenses and give rise to the charge order in (f). The data that determine of the transition points are given in Supplemental Material (SM) [50].
In the past, the dynamic (spectral) properties of FCI are mainly studied in projected topological flat band, mimic the investigations of the charge neutral excititaions with Landau level projection in FQHE [51, 39, 33, 34, 37, 41, 45, 35, 52, 53]. While the universal properties could be undoubtly studied in this way, the dynamics are model dependent and largely affected by band mixing, which could destroy FCI by quantum fluctuation and might be responsible for the sharp difference of the particle-hole fillings of the same Chern band in experiments [54, 55, 13]. Moreover, previous studies of the dynamic properties of FCI focused on the charge excitations [56], which is typically of a higher energy scale compared with neutral excitations. Although the single mode approximation (SMA) was done for FCI (projecting to one Chern band and no spectrum weight) [52], the low-lying dispersions with the identification of roton minimum were not clear without mapping the quantum numbers of momenta between discrete lattices and continuum models, which was model-dependent, and the importance of the spectrum weight has not been revealed. While the energy scale of different excitations in FCI has been discussed through thermodynamic calculation [25, 26], a comprehensive study of the dynamic properties of FCI is still lacking. However, it is well understood that such direct and spectroscopic evidence for the presence of magnetorotons and chiral gravitons, will be of great importance to guide future experiments and provide unbiased verifications for the developments of unifying theories for FQHE and FCI systems.
@@
In this letter, we fill in such a knowledge gap by studying the spectral properties of FCI realised by hard-core bosons on Haldane honeycomb lattice model at half filling of the flat band [5, 4]. And the FCI therein can transit to two CDW phases as a function of the repulsive density-density interaction [29]. We employ the state-of-the-art time-dependent variation principle (TDVP) [57, 58] within large-scale density matrix renormalization group (DMRG) simulaltion [59] to reveal the spectroscopic signatures for the FCI state and its transition to CDW orders. We find that (1) the charge neutral magnetoroton can be visualized from the dynamic density structure factor. The roton mode could go soft and finally leads to the CDW order of the same momentum as the roton minimum, whereas the single-particle gap remains large and intact throughout; (2) The spectral weight of roton become sharper and highly concentrated around its minimum when approaching the phase boundary, manifesting the roton condensation character of the FCI-CDW transition; (3) Inside the FCI, a strong gapped chiral graviton mode – probed by quadrupolar density correlation function – with angular momentum-2 is clearly present and its chirality can be tuned by flipping the Chern number ±$\pm 1$ of the flat band. Our results shows, in an unbiased manner, the robustness and tunability of the charge neutral excitations in the FCI states and provide guidance towards their spectroscopic detection in quantum moiré materials.
Model and Method.— We consider the bosonic topological flat band Hamiltonian on honeycomb lattice [5, 4, 29]:
@@
In our simulation, we consider a finite cylinder with $N_{1}\times N_{2}=4\times 16$ unit cells. The total number of sites is $N=2N_{1}N_{2}$ corresponding bosons particle number of $1/2$ fills the flat band is $N_{b}=N_{1}N_{2}/2=32$. The details of the cylinder geometry and simulations is described in SM [50].
Before discussing the spectra, we begin with the ground state phase diagram as shown in Fig.1, The system go through a FCI-Solid I-Solid II phase transition triggered by the condensation of two successive rotons (with momenta $\mathbf{b_{1}}/2$ and $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ respectively) as we increase $V_{2}$ along the path $V_{1}=4$. Such successive phase transition can be interpreted as a vestigial transition [60, 61, 62, 63, 64, 65] with the translational symmetry of the lattice broken in a two-step manner [29]. Here, static density structure factor $S^{A}(\mathbf{k})=\frac{1}{N_{A}}\sum_{ij}e^{-i\mathbf{k}\cdot(\mathbf{x_{i}-x_{j}})}\langle n_{i,A}n_{j,A}\rangle$ and charge order parameter $\delta^{A}(\mathbf{k})=\frac{1}{\sqrt{N_{A}}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{x_{i}}}\langle n_{i,A}\rangle$ are measured to diagnosis the phases, where $A$ denotes the sub-lattice, $N_{A}$ denotes its site number, $i,j$ run over all unit cells. Since the results are the same for both sublattices, we mainly present the results on $A$ sublattice. The bright point at $\frac{\mathbf{b_{1}}}{2}$ in Fig. 1 (a) denotes the location of the lowest roton minimal of FCI [25, 26, 29], the fact that the rotational symmetry is not recovered in the $S^{A}(\mathbf{k})$ is a consequence of the cylindrical geometry of the DMRG simulation. In this system there is a second roton in the FCI phase, and it is located at the $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ (it can be clearly visualised in the spectra in Fig. 2 (b)), and once the lowest roton condenses and the system enters the Solid I phase, the measurement of $S^{(A)}(\mathbf{k})$ in Fig. 1 (b) shows that the brightest density correlation is at this wave vector ($V_{1}=4,V_{2}=3.2$). The charge order of Solid I, however, can be seen from its order parameter $\delta^{A}(\mathbf{k})$ in Fig. 1 (e) ($V_{1}=4,V_{2}=3.2$). And the grey arrow connecting the momentum $\mathbf{b_{1}}/2$ from Fig. 1 (a) to Fig. 1 (e), symbolises the condensation of the roton and formation of the charge order. The process also happened in transition from Solid I to Solid II, symbolised by the grey arrow connecting the momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ of the second roton in Fig. 1 (b) to the charge order in Fig. 1 (f) ($V_{1}=4,V_{2}=4.6$).
@@
where $n_{j,A}(t_{l})=e^{iHl\Delta t}n_{j,A}e^{-iHl\Delta t}$ as $l$ runs from 00 to $N_{t}$ and $\eta=0.05$. We time evolve $n_{0,A}|\psi_{0}\rangle$ with $|\psi_{0}\rangle$ the ground state wave function and $n_{0,A}$ defined in the bulk of the cylinder [50]. The correlation function $\langle\hat{n}_{j,A}(t)\hat{n}_{0,A}\rangle$ is measured at every time slice as we progressively evolve $n_{0,A}|\psi_{0}\rangle$ with operator $e^{-iHt_{l}}$ in TDVP. For $N_{t}$, $n_{0,A}$ and the details of spectrum calculation, see SM [50].
@@
Figure 2: Magnetoroton spectrum inside FCI phase. (a), (b) are $S^{A}(\mathbf{k},\omega)$ deep inside FCI with $V_{1}=4,V_{2}=1$ along Path I and Path III in Fig. 1 (d). Two charge neutral rotons relevent with the CDW phases are marked with $\Delta_{n_{1}}$ and $\Delta_{n_{2}}$ with their gaps denoted by the red and yellow arrows, respectively. The dashed lines trace the dispersion relation of roton. (c) and (d) are the roton spectra near the FCI-Solid I phase boundary with $V_{1}=4,V_{2}=2.5$. The $\Delta_{n_{1}}$ is about to close with very sharp spectral weight and give rise to the Solid I phase with the charge order at same momentum. The $\Delta_{n_{2}}$ will persist into the Solid I phase and condense at the Solid I - Solid II transition (discussed in Fig. 1 (b) and (f)).
@@
We also measure the single-particle Green’s functions $\langle b_{j,A}(t_{l})b^{\dagger}_{0,A}\rangle$ and $\langle b^{\dagger}_{j,A}(t_{l})b_{0,A}\rangle$ respectively and label the corresponding spectra with $G_{e}^{A}(\mathbf{k},\omega)$ (electron) and $G_{h}^{A}(\mathbf{k},\omega)$ (hole). We sum $G_{e}^{A}(\mathbf{k},\omega)$ and $G_{h}^{A}(\mathbf{-k},-\omega)$ to obtain single-particle spectrum $G^{A}(\mathbf{k},\omega)$, where we use the particle-hole conjugation relation to reverse the sign of momentum and energy for holes. Again $b^{\dagger}$ and $b$ locate at the bulk of the cylinder. The local density of state can then be obtained as $\rho(\omega)=\frac{1}{N_{k}}\sum_{\mathbf{k}}G^{A}(\mathbf{k},\omega)$.
@@
Fig. 2 shows the $S^{A}(\mathbf{k},\omega)$ along the path I and III in the BZ in Fig. 1 (d). Panels (a) and (b) of Fig. 2 are deep inside the FCI phase, and one clearly sees the lowest roton minimal at $M_{1}$ with the roton gap $\Delta_{n_{1}}\sim 1$, which is much smaller the single-particle gap computed at the same parameter which is $\Delta_{c}\sim 2$ with sharper spectrum weight, as shown in Fig. 3 (a). This observation further reinforces the understanding that in the FCI phase, the lowest excitations are the charge neutral roton instead of the single-particle ones [25]. Panel (b) exhibits the similar lowest rotons at $M_{3}$ and $M_{5}$, but more interestingly, there is another higher energy roton manifesting at momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$, this is the second roton that will trigger the vestigial transition from Solid I to Solid II, once the lowest roton $\Delta_{n_{1}}$ condense to give rise to the Solid I. In Fig. 2 (c) and (d), the parameter $(V_{1}=4,V_{2}=2.5)$ is closer to the FCI-Solid I transition, and one sees that the roton gap $\Delta_{n_{1}}$ becomes smaller, signifying the approaching transition.
@@
We note that as the roton is about to condense, its spectral weight become much sharper and concentrated close to its minimum compared with that deep inside the FCI state, we expect such clear dynamic response of the magnetoroton of FCI close to its transition to CDW phase, can be seen in the future scanning tunneling spectrospcopy (STS) or nano-ARPES experiments. Moreover, our previous thermodynamic calculation showed that, in the correlated flat band systems, close to the energy scale of the neutral gap, the compressibility of the system is greatly enhanced [66, 25, 27, 29, 67], here one sees similar proliferation of the roton excitation happen at the FCI-CDW transition and expects there will be strong signal in the compressibility which can be probed from the quantum capacitance experiment.
@@
Figure 3: Single-particle spectrum inside the FCI phase. (a), (b) are the $G^{A}(\mathbf{k},\omega)$ and $\rho(\omega)$ inside the FCI phase (with the same parameter as Fig. 2 (a) and (b)) and close to the FCI-Solid I transition (with the same parameter as Fig. 2 (c) and (d)), respectively. In both cases, the single-particle gap $\Delta_{c}\approx 2$ are much larger than the roton gaps $\Delta_{n_{1}}$ and $\Delta_{n_{2}}$. The path (Path I) we have chosen to display the single-particle spectrum features a minimal charge gap and encompasses major of the total spectral weight.
@@
We also monitor the evolution of the single-particle spectra along the same parameter path, and the results are shown in Fig. 3. We plot $G^{A}(\mathbf{k},\omega)$ and $\rho(\omega)$ in Fig. 3 (a) with $V_{2}=1$ (well inside the FCI phase) and Fig. 3 (b) with $V_{2}=2.5$ (close to the FCI-Solid I transition). It is clear that at both parameters, the single-particle gaps are well developed and much larger than the roton gaps at the same parameter. Such results suggest that, the transition out of a FCI can be purely driven by the charge-neutral roton modes, without the closing of the single-particle gap [26, 27, 28, 29].
@@
Chiral Gravitons.— The operators for measuring CGM in lattice are given by the representation of $d$-wave operators with angular momentum-2 along the NN bonds, $d_{x^{2}-y^{2}}=\sum_{i}n_{r_{i}}n_{{r_{i}}+\bm{a}_{1}}-\frac{1}{2}n_{r_{i}}n_{{r_{i}}+\bm{a}_{2}}-\frac{1}{2}n_{r_{i}}n_{{r_{i}}+\bm{a}_{3}}$ and $d_{xy}=\sum_{i}\frac{\sqrt{3}}{2}\left(n_{r_{i}}n_{{r_{i}}+\bm{a}_{2}}-n_{i}n_{{r_{i}}+\bm{a}_{3}}\right)$, where the form factor depends on the unit vector $\bm{a}_{i=1,2,3}$ as depicted in Fig. 1 (c). The chiral density operator $O^{\pm}=d_{x^{2}-y^{2}}\pm id_{xy}$ is constructed from the $d$-wave operators that is angular momentum $S=2$ or $S=-2$ under chiral transformation [41, 40]. We therefore define the dynamic quadrupolar density correlation function as
$g^{\pm}(\mathbf{k},\omega)$ $=\frac{1}{\sqrt{N_{t}N_{A}}}\sum_{jl}e^{i\omega t_{l}}e^{-i\mathbf{k}\cdot(\mathbf{x_{j}}-\mathbf{x_{0}})}\langle(O^{\pm}_{j,A}(t_{l}))^{\dagger}O^{\pm}_{0,A}\rangle$ (3)
$-\langle(O^{\pm}_{j,A})^{\dagger}\rangle\langle O^{\pm}_{0,A}\rangle$
with $(O^{\pm})^{\dagger}=O^{\mp}$.

### v3

@@ Abstract
Employing the state-of-the-art time-dependent variational principle (TDVP) algorithm, we compute the spectra of charge-neutral excitations in the $\nu=1/2$ (bosonic) and $1/3$ (fermionic) fractional Chern insulator (FCI) on the Haldane honeycomb lattice model. The magnetoroton visualized from the dynamic density structure factor acquires a minimum gap at finite momentum that can go soft with increasing interaction and give rise to a charge density wave (CDW) at the same wavevector. As the system approaches the FCI-to-CDW transition point, we observe a pronounced sharpening of the roton mode, suggesting that the magnetoroton behaves more like a quasiparticle as it softens. Notably, this occurs while the single-particle gap remains finite. Besides the magnetoroton at finite momentum, we also construct quadrupolar chiral operators in a discrete lattice and resolve the chiral graviton mode around the $\Gamma$ point of the Brillouin zone. Furthermore, we show the different chiralities of the gravitons of FCIs with opposite-sign Hall conductance for the first time. Our results offer the clear spectral observations of magnetoroton and chiral graviton in FCI lattice models and will have relevance towards systematic understanding of elementary excitation in FCIs.
Introduction.— Charge-neutral excitation in fractional quantum Hall (FQH) states has drawn immense attention in the past decades. The pioneer work by Girvin, MacDonald, and Platzman (GMP) established the magnetoroton theory of the Laughlin wave function based on the single-mode approximation (SMA) [1], which interpreted the charge-neutral mode as a collective density excitation [2, 3, 4, 5, 6] whose softening triggers a phase transition from FQH to CDW [1, 7, 8, 9]. Later on, Haldane propose a geometric description of charge neutral mode on long wavelength limit, reflecting oscillation of intrinsic metric of Landau orbitals [10], known as chiral graviton mode (CGM) with spin angular momentum $2$  [11, 12, 13, 14, 15, 16, 17]. CGM has recently been observed in the experiment using circularly polarized resonant inelastic light scattering [18]. While the magnetoroton mode tends to break translational symmetry and trigger FCI-CDW transition, the graviton mode could break rotational symmetry and is responsible for anisotropic transport in the FQH nematics  [19, 20, 21, 22] and describes nonequilibrium dynamics after magnetic quench [23, 24].
FCI is the lattice analog of FQH [25, 26, 27, 28] and could exist without external magnetic field due to the orbital magnetism [26, 29, 30]m, also known as the fractional quantum anomalous Hall (FQAH) states. FQAH was recently observed in MoTe2 bilayers [31, 32, 33, 34] and in rhombohedral pentalayer graphene/hBN moiré superlattices [35]. In parallel with FQH, the study of charge-neutral excitations in FCI mostly adopts GMP ansatz, where the Hilbert space is restricted to a single band and perform exact diagonalization (ED), yielding similar behavior of charge neutral excitation as FQH [36, 14, 1, 7, 12, 16, 20, 17, 37, 38, 39]. However, the investigation of FCI present several unique aspects. The energy scales in the FCI system are more complex, where the interaction strength could be larger than the band gap, making the band mixing effect non-negligible [40, 41, 35]. Moreover, the importance of spectral weight remains unrevealed, as it contains information on quasi-particle lifetime and is directly related to full width at half maximum analysis in the Raman scattering experiment [18]. Furthermore, investigations of CGMs have predominantly employed momentum-space approaches, while the detection of CGM based on a real-space approach (considering the lattice effect) remains largely unexplored. Given the situation, a comprehensive study of the spectral properties of magnetoroton and CGM in FCI will be of great importance to guide future experiments and provide unbiased verifications for the development of unifying theories for FQH and FCI systems.
@@
Figure 1: Static density structure factor $S^{A}(\mathbf{k})$ and charge order parameter $\delta^{A}(\mathbf{k})$ for $1/2$ bosonic FCI. Along the parameter path we choose, the system undergoes a sequence of FCI-Solid I-Solid II phase transitions. (a), (b) are the $S^{A}(\mathbf{k})$ of FCI phase ($V_{1}=4,V_{2}=2$) and Solid I phase ($V_{1}=4,V_{2}=3$). (c), (d) are the real and reciprocal spaces of the honeycomb lattice FCI model in Eq. (1). In (c), $\mathbf{r_{1}}=(\sqrt{3},0),\mathbf{r_{2}}=(-\sqrt{3}/2,3/2)$ are the primitive lattice vectors. The NN, NNN, and NNNN hoppings are depicted with black, blue and green colours, respectively. In (d), the solid hexagon is the first Brillouin zone (BZ) of the model, with high symmetry points labelled. The four paths that are discrete in $\mathbf{b_{2}}$ direction with step $\mathbf{b_{2}}/4$ are where we compute the spectra under the cylinder geometry in DMRG calculation. (e), (f) are the $\delta^{A}(\mathbf{k})$ of Solid I phase ($V_{1}=4,V_{2}=3$) and Solid II phase ($V_{1}=4,V_{2}=5$). The roton mode locates at $\mathbf{b_{1}}/2$ in (a) condensates and leads to charge order in (e), while similar $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ roton in (b) condenses and give rise to the charge order in (f). The data that determine of the transition points are given in Supplemental Material (SM) [42].
@@
In this letter, we fill in such a knowledge gap by studying the spectral properties of FCI realised by hard-core bosons and spinless fermions on the Haldane honeycomb lattice model when the flat band is partially filled [27, 26]. And the $1/2$ bosonic FCI therein can transit to two CDW phases as a function of the repulsive density-density interaction [43]. We employ the state-of-the-art time-dependent variation principle (TDVP) [44, 45] within large-scale density matrix renormalization group (DMRG) simulaltion [46] to reveal the spectroscopic signatures for the FCI state and its transition to CDW orders. We find that (1) the charge-neutral magnetoroton can be visualized from the dynamic density structure factor. The magnetoroton mode could go soft and finally leads to the CDW order at the same momentum of the roton minimum, whereas the single-particle gap remains large and intact throughout; (2) the spectral weight of roton become sharper and highly concentrated around its minimum when approaching the phase boundary, manifesting the roton condensation character of the FCI-CDW transition; (3) inside the $1/2$ bosonic FCI, a strong gapped chiral graviton mode – probed by quadrupolar density correlation function – with angular momentum-2 is clearly present and its chirality can be tuned by flipping the Chern number ±$\pm 1$ of the flat band; (4) the chiral signal observed in the $1/3$ fermionic FCI exhibits significantly reduced spectral weight compared to the bosonic case, suggesting potentially distinct behavior of CGM in fermionic FCI. Our results show, in an unbiased manner, the robustness and tunability of the charge-neutral excitations in the FCI states and provide guidance towards their spectroscopic detection in experiments.
Model and Method.— We consider the topological flat band Hamiltonian on honeycomb lattice [27, 26, 47, 43, 48] (in the analysis of magnetoroton and charge excitation in the text, we focus on the $1/2$ bosonic FCI, which could be realized in cold atom system [49], leaving the $1/3$ fermionic FCI results in the supplemental material(SM) [42]. While for CGM in the main text we display both $1/2$ bosonic FCI and $1/3$ fermionic FCI results):
@@
In our simulation, we consider a finite cylinder with $N_{1}\times N_{2}=4\times 16$ unit cells. The total number of sites is $N=2N_{1}N_{2}$ corresponding bosons particle number of $1/2$ fills the flat band is $N_{b}=N_{1}N_{2}/2=32$. For $1/3$ fermionic FCI, we consider a $N_{1}\times N_{2}=3\times 18$ cylinder. The details of the cylinder geometry and simulations are described in SM [42].
Before discussing the spectra, we begin with the ground state phase diagram as shown in Fig.1, The system goes through a FCI-Solid I-Solid II phase transition triggered by the condensation of two successive rotons (with momenta $\mathbf{b_{1}}/2$ and $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ respectively) as we increase $V_{2}$ with fixed $V_{1}=4$. Such a successive phase transition is similar with a vestigial transition [50, 51, 52, 53, 54, 55] with the translational symmetry of the lattice broken in a two-step manner [43]. Here, static density structure factor $S^{A}(\mathbf{k})=\frac{1}{N_{A}}\sum_{ij}e^{-i\mathbf{k}\cdot(\mathbf{x_{i}-x_{j}})}\langle n_{i,A}n_{j,A}\rangle$ and charge order parameter $\delta^{A}(\mathbf{k})=\frac{1}{\sqrt{N_{A}}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{x_{i}}}\langle n_{i,A}\rangle$ are measured to diagnosis the phases, where $A$ denotes the sub-lattice, $N_{A}$ denotes its site number, $i,j$ run over all unit cells. Since the results are the same for both sublattices, we mainly present the results on $A$ sublattice. The bright point at $\frac{\mathbf{b_{1}}}{2}$ in Fig. 1 (a) denotes the location of the lowest roton minimal of FCI [56, 57, 43], the fact that the rotational symmetry is not recovered in the $S^{A}(\mathbf{k})$ is a consequence of the cylindrical geometry of the DMRG simulation. In this system there is a second roton in the FCI phase, and it is located at the $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ (it can be clearly visualised in the spectra in Fig. 2 (b)), and once the lowest roton condenses and the system enters the Solid I phase, the measurement of $S^{(A)}(\mathbf{k})$ in Fig. 1 (b) shows that the brightest density correlation is at this wave vector ($V_{1}=4,V_{2}=3$). The charge order of Solid I, however, can be seen from its order parameter $\delta^{A}(\mathbf{k})$ in Fig. 1 (e) ($V_{1}=4,V_{2}=3$). And the grey arrow connecting the momentum $\mathbf{b_{1}}/2$ from Fig. 1 (a) to Fig. 1 (e), symbolises the condensation of the roton and formation of the charge order. The process also happened in transition from Solid I to Solid II, symbolised by the grey arrow connecting the momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ of the second roton in Fig. 1 (b) to the charge order in Fig. 1 (f) ($V_{1}=4,V_{2}=5$). The real space charge occupation pattern of two Solid phases is shown in SM [42].
@@
where $n_{j,A}(t_{l})=e^{iHl\Delta t}n_{j,A}e^{-iHl\Delta t}$ as $l$ runs from 00 to $N_{t}$ and we choose $\eta=0.05$ to broaden the peaks. We time evolve $n_{0,A}|\psi_{0}\rangle$ with $|\psi_{0}\rangle$ the ground state wave function and $n_{0,A}$ defined in the bulk of the cylinder [42]. The correlation function $\langle\hat{n}_{j,A}(t)\hat{n}_{0,A}\rangle$ is measured at every time slice as we progressively evolve $n_{0,A}|\psi_{0}\rangle$ with operator $e^{-iHt_{l}}$ in TDVP. For $N_{t}$, $n_{0,A}$ and the details of spectrum calculation, see SM [42].
@@
Figure 2: Magnetoroton spectrum inside $1/2$ bosonic FCI phase. (a), (b) are $S^{A}(\mathbf{k},\omega)$ deep inside FCI with $V_{1}=4,V_{2}=1$ along Path I and Path III in Fig. 1 (d). Two charge neutral rotons relevant with the CDW phases are marked with $\Delta_{n_{1}}$ and $\Delta_{n_{2}}$ with their gaps denoted by the red and yellow arrows, respectively. (c) and (d) are the roton spectra near the FCI-Solid I phase boundary with $V_{1}=4,V_{2}=2$. The $\Delta_{n_{1}}$ is about to close with very sharp spectral weight and give rise to the Solid I phase with the charge order at same momentum. The $\Delta_{n_{2}}$ will persist into the Solid I phase and condense at the Solid I - Solid II transition (discussed in Fig. 1 (b) and (f)).
@@
We also measure the single-particle Green’s functions $\langle b_{j,A}(t_{l})b^{\dagger}_{0,A}\rangle$ and $\langle b^{\dagger}_{j,A}(t_{l})b_{0,A}\rangle$ respectively and label the corresponding spectra with $G_{e}^{A}(\mathbf{k},\omega)$ (electron) and $G_{h}^{A}(\mathbf{k},\omega)$ (hole). We sum $G_{e}^{A}(\mathbf{k},\omega)$ and $G_{h}^{A}(\mathbf{-k},-\omega)$ to obtain single-particle spectrum $G^{A}(\mathbf{k},\omega)$, where we use the particle-hole conjugation relation to reverse the sign of momentum and energy for holes. $b^{\dagger}$ and $b$ locate at the bulk of the cylinder. The local density of state can then be obtained as $\rho(\omega)=\frac{1}{N_{k}}\sum_{\mathbf{k}}G^{A}(\mathbf{k},\omega)$.
@@
Fig. 2 shows the $S^{A}(\mathbf{k},\omega)$ along the path I and III in the BZ in Fig. 1 (d). Panels (a) and (b) of Fig. 2 are deep inside the FCI phase, and one clearly sees the lowest roton minimal at $M_{1}$ with the roton gap $\Delta_{n_{1}}\sim 1$, which is much smaller the single-particle gap computed at the same parameter which is $\Delta_{c}\sim 7$ with sharper spectrum weight, as shown in Fig. 3 (a). This observation further reinforces the understanding that in the FCI phase, the lowest excitations are the charge neutral roton instead of the single-particle ones [56, 58, 59]. Panel (b) exhibits the similar lowest rotons at $M_{3}$ and $M_{5}$, but more interestingly, there is another higher energy roton manifesting at momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$, this is the second roton that will trigger the vestigial transition from Solid I to Solid II, once the lowest roton $\Delta_{n_{1}}$ condense to give rise to the Solid I. In Fig. 2 (c) and (d), the parameter $(V_{1}=4,V_{2}=2)$ is closer to the FCI-Solid I transition, and one sees that the roton gap $\Delta_{n_{1}}$ becomes smaller, signifying the approaching transition. For the discussion of magnetoroton in $1/3$ fermionic FCI, please refer to SM [42].
@@
We note that as the roton is about to condense, its spectral weight become much sharper and concentrated close to its minimum compared with that deep inside the FCI state, we expect such clear dynamic response of the magnetoroton of FCI close to its transition to CDW phase, can be seen in the future Raman scattering experiments. Moreover, our previous finite-temperature calculations have shown the thermodynamic signals of the neutral excitations of (F)QAH states in the correlated flat band systems. As the temperature approaches the energy scale of the neutral gap, the compressibility of the system is greatly enhanced [60, 56, 59, 43, 61, 62],.
@@
Figure 3: Single-particle spectrum inside the FCI phase. (a), (b) are the $G^{A}(\mathbf{k},\omega)$ and $\rho(\omega)$ inside the FCI phase (with the same parameter as Fig. 2 (a) and (b)) and close to the FCI-Solid I transition (with the same parameter as Fig. 2 (c) and (d)), respectively. In both cases, the single-particle gap $\Delta_{c}\approx 7$ are much larger than the roton gaps $\Delta_{n_{1}}$ and $\Delta_{n_{2}}$. The path (Path I) we have chosen to display the single-particle spectrum features a minimal charge gap and encompasses major of the total spectral weight.
@@
We also monitor the evolution of the single-particle spectra along the same parameter path, and the results are shown in Fig. 3. We plot $G^{A}(\mathbf{k},\omega)$ and $\rho(\omega)$ in Fig. 3 (a) with $V_{2}=1$ (well inside the FCI phase) and Fig. 3 (b) with $V_{2}=2$ (close to the FCI-Solid I transition). It is clear that at both parameters, the single-particle gaps are well developed and much larger than the roton gaps at the same parameter. Such results suggest that, the transition out of a FCI can be purely driven by the charge-neutral roton modes, not related the single-particle gap [57, 59, 58, 43].
@@
Chiral Gravitons.— We combine the quadrupolar density operator with chiral form factor to probe CGM phenomenologically. We define $O^{\pm}_{m}=\sum_{l}e^{\pm iL_{z}\phi_{l}}n_{r_{m}}n_{{r_{m+\delta_{l}}}}$, where $l\in\{1,2,\dots,6\}$ connect site $m$ with it’s 6 next nearest neighbor on same sub-lattice and $\phi_{l}=l\frac{\pi}{3}$ with the $l$ runs counterclockwise [16, 15], such that $O^{\pm}_{m}$ create a chiral excitation with angular momentum ±$\pm 2$ at site $m$. We therefore define the correlation function of $O^{\pm}_{m}$
$G^{\pm}(\mathbf{k},\omega)$ $=\frac{1}{\sqrt{N_{t}N_{A}}}\sum_{jl}e^{i(\omega+i\eta)t_{l}}e^{-i\mathbf{k}\cdot(\mathbf{x_{j}}-\mathbf{x_{0}})}\langle O^{\mp}_{j}(t_{l})O^{\pm}_{0}\rangle$ (3)
with $\eta=0.05$, $(O^{\pm})^{\dagger}=O^{\mp}$, In the calculation of Eq. (3), the summation over the sites runs over the same sublattice. The chirality of the CGM is determined by comparing the spectrum weight of $G^{\pm}(\mathbf{k},\omega)$ at $\Gamma$ point with $G^{+}(\mathbf{k},\omega)$ probes $S=-2$ signal vise versa). We further combine $g^{+}(\mathbf{k},\omega)=G^{+}(\mathbf{k},\omega)-G^{-}(\mathbf{k},\omega)$ and $g^{-}(\mathbf{k},\omega)=G^{-}(\mathbf{k},\omega)-G^{+}(\mathbf{k},\omega)$ such that their subtraction cancels the non-chiral part of the spectral signature. We note that $g^{\pm}(\mathbf{k},\omega)$ is also the time-ordered Green’s function of $O^{\pm}_{m}$.